four years hence a boy will be twice as old as he was 6 years ago.find his present age?

Answer by checkley79(3341)

X+4=2(X-6)
X+4=2X-12
X-2X=-12-4
-X=-16
X=16 AGE NOW.
PROOF:
16+4=2(16-6)
20=2*10
20=20